OpenStudy (anonymous):

ALGEBRA HELP PLEASE? WILL GIVE MEDAL + FAN!!! Melania has two jobs. During the day, she works as an office clerk, and in the evening she works as a cashier. Her office job pays her $13 per hour, and her cashier job pays her $9.25 per hour. In one week, Melania worked 38 hours, and she earned a total of $434. How much did she earn that week working at her cashier job?

OpenStudy (leader):

so her office job is o and her cashier job is c

OpenStudy (leader):

Then o +c =38

OpenStudy (leader):

And the second equation is 13o+ 9.25c= 434

OpenStudy (leader):

they intersect at (22,16) where she worked 16 hour as a cashier.

OpenStudy (leader):

So 16 * 9.25 = $148

OpenStudy (leader):

So 148 is your final answer.

OpenStudy (anonymous):

That was very simple thanks :D I'll fan u in a sec

OpenStudy (leader):

Thanks :)